OpenStudy (anonymous):

medal for correct answer!!!!!!!! which set of points can be modeled without a power model? a) {(2.2, 4.33), (3.5, 6.641), (4.8, 10.184)} b) {(1.4, 0.133), (2, 0.21), (2.6, 0.333)} c) {(1.1, 17.155), (1.7, 14.941), (2.3, 13.013)} d) {(1.4, 5.192), (1.8, 6.104), (2.2, 7.016)}

OpenStudy (anonymous):

Examine the slopes {(1.4, 5.192),(1.8, 6.104),(2.2, 7.016)} point 1 and point 2: (6.104-5.192)/(1.8-1.4) = 2.28 (7.016-6.104)/(2.2-1.8) = 2.28 equal slopes can be modeled without a power model.